Let X be a binomial(5, 0.2) random variable. Let Y be a discrete random variable that is independent of X, such that Y = 1 with probability 0.2 and Y = 0 with probability 0.8. What is the probability that the sum of X and Y is less than or equal to 3?


1
Expert's answer
2022-03-15T19:38:15-0400

"P(X=x)=(\begin{matrix}
 5 \\
 x
\end{matrix})0.2^x(1-0.2)^{5-x}"




"P(Y=0)=0.8,P(Y=1)=0.2"

"P(X+Y\u22643)=P(Y=0,X\u22643)+P(Y=1,X\u22642)"

"=0.8P(X\u22643)+0.2P(X\u22642)"

"=0.8(0.32768+0.4096+0.2048+0.0512)+0.2(0.32768+0.4096+0.2048)"

"=0.98304"